# A Purely Data-Driven Adaptive Impedance Matching Method Robust to Parasitic Effects

### v1

@@ Abstract
Adaptive impedance matching between antennas and radio frequency front-end (RFFE) power modules is essential for mobile communication systems. To address the matching performance degradation caused by parasitic effects in practical tunable matching networks (TMN), this paper proposes a purely data-driven adaptive impedance matching method that avoids trial-and-error physical adjustment. First, we propose the residual enhanced circuit behavior modeling network (RECBM-Net), a deep learning model that maps TMN operating states to their scattering parameters (S-parameters). Then, we formulate the matching process based on the trained surrogate model as a mathematical optimization problem. We employ two classic numerical methods with different online computational overhead, namely simulated annealing particle swarm optimization (SAPSO) and adaptive moment estimation with automatic differentiation (AD-Adam), to search for the matching solution. To further reduce the online inference overhead caused by repeated forward propagation through RECBM-Net, we train an inverse mapping solver network (IMS-Net) to directly predict the optimal solution. Simulation results show that RECBM-Net achieves exceptionally high modeling accuracy. While AD-Adam significantly reduces computational overhead compared to SAPSO, it sacrifices slight accuracy. IMS-Net offers the lowest online overhead while maintaining excellent matching accuracy.
@@ I Introduction
The impedance mismatch between the antenna and the radio frequency (RF) front-end power module reduces the power transmitted to the antenna and degrades overall system performance in mobile communication systems. Furthermore, impedance mismatch adversely affects the linearity of the digital pre-distortion power amplifier [1], resulting in signal modulation distortion and spectrum expansion. More critically, the reflected power caused by the mismatch may damage the RF front-end (RFFE) components [2]. Therefore, achieving impedance matching is crucial for mobile devices to maintain reliable and high-performance communication.
@@
The antenna impedance of mobile devices is influenced by multiple factors, e.g. operating frequency [3], various electromagnetic interactions arising from user holding methods [4, 5], user proximity [6, 7], and even user age and clothing [8], etc. The dynamic operating conditions of these devices induce impedance mismatches in the time-frequency domain. To mitigate the detrimental effects of impedance mismatch, extensive research has been conducted on adaptive impedance matching techniques for mobile communication systems.
@@
Traditional adaptive impedance matching is primarily achieved through numerical methods. The numerical methods use the feedback signal reflecting the degree of mismatch to determine the optimal matching solution through extensive trial and error. The works in [9, 10, 11] adopted sequential search strategies based on matching performance feedback. To enhance search efficiency, the authors in [12] employed various gradient-free optimization methods to minimize the magnitude of the reflection coefficient, such as the single-step and powell algorithms. Additionally, the work in [13, 14, 15] employed heuristic algorithms, while the work in [16] utilized gradient information to perform steepest gradient descent, reducing the number of trail-and-error iterations. The authors in [17] observed that when the parallel inductance of the T-network is fixed, the real part of the input impedance follows a linear fractional relationship with the reactance of the first series element, and proposed a binary search algorithm to reduce the number of iterations. The disadvantages of numerical methods are their inefficiency and riskiness. As the matching network components must be physically adjusted during each trial, this tuning process introduces both amplitude and phase modulation to the radiated signal, which can lead to data corruption if tuning occurs during transmission [18]. Therefore, numerical methods prove impractical for achieving adaptive impedance matching in mobile devices.
@@
Compared to numerical methods, analytical methods derive the matching solution directly from accurate complex impedance information and the circuit model, thereby avoiding the iterative trial and error procedures. The authors in [19] proposed a method for real-time measurement of the complex impedance at any location within the RFFE of a handset based on perturbation theory, providing critical support for the analytical computation of the matching solution. The work of [20] presented explicit mathematical expressions for the values of components when the ideal $\pi$-network is perfectly matched. Furthermore, the authors in [21] developed an orthogonal detection technique that directly sets the real and imaginary parts of the impedance. To address imperfect matching caused by the limited tuning range of practical $\pi$-network, the work of [22] proposed an analytical method to compute the optimal solution within the tunable range. Additionally, the work of [23] considered the parasitic effects introduced by MEMS components in the practical $\pi$-network and provided an analytical solution by solving the conjugate matching equation of the equivalent circuit, which remains an approximation of the practical circuit. As analytical methods are inherently model-dependent, the accuracy of the matching is determined by the degree of conformity between the practical circuit model and the model assumed by the algorithm.
@@
Recently, powerful artificial intelligence (AI) technologies have been applied to achieve efficient and accurate adaptive impedance matching. To address impedance mismatch in the frequency domain, the work in [24] proposed a deep learning model that maps the desired reflection coefficient frequency response to the corresponding L-network configuration for a given planner inverted-F antenna. Considering the complexity of the model, the authors in [25] proposed a low-complexity shallow learning model based on ridge regression to achieve the same mapping. In addition, to address the impedance mismatch caused by variations in the distance between the transmitting and receiving coils in wireless power transfer systems, the work in [26] employed a neural network to establish a mapping between the equivalent load impedance and the corresponding matching solution at a specific frequency. Furthermore, considering time-frequency domain mismatch, the work of [27] proposed a deep neural network (DNN) to directly determine the ideal $\pi$-network matching solution based on incomplete impedance information. The construction of large and accurate supervised datasets is critical to the success of AI-based adaptive impedance matching methods.
@@
The above analytical and AI-based methods are applicable only to tunable matching networks (TMN) that either neglect parasitic effects or consider only simplified cases. However, parasitic effects in practical tunable matching networks (TMN) are unavoidable and complex due to inherent limitations in the process of manufacturing and packaging of RF circuits. For example, common structures such as traces, bonding wires [28], vias [29], pins [30], solder balls [31], etc., will introduce parasitic capacitance, inductance, and resistance. Particularly in mobile communication systems, TMNs are often implemented using discrete electronically switchable inductor and capacitor banks [32] to satisfy miniaturization constraints. The dense integration of switches, transmission lines, and non-ideal discrete components significantly aggravates parasitic effects.
@@
Parasitic effects render the equivalent circuit topology of a practical TMN unknown. When these effects are taken into account, matching solutions derived from analytical methods based on idealized or specific topologies often fail to meet the actual matching requirements. Similarly, for the aforementioned AI-based methods, datasets constructed from idealized topologies fail to reflect real matching behavior, while those based on practical topologies suffer from the difficulty of acquiring accurate labels. These issues lead to degraded matching performance in AI-based methods as well. To the best of our knowledge, achieving efficient and accurate adaptive impedance matching remains an open problem when parasitic effects cause the practical TMN to deviate from its idealized topology.
@@
To address this challenge, in this paper, we propose a purely data-driven adaptive impedance matching method. First, we analyze the impact of parasitic effects on the behavior of TMN using multiport network theory. Considering the high nonlinearity and uncertainty introduced by these parasitic effects, we propose a deep learning model, residual enhanced circuit behavior modeling network (RECBM-Net), to accurately learn the system behavior. Based on the trained RECBM-Net, we further develop three matching solution determination strategies with varying online computational overhead: simulated annealing particle swarm optimization (SAPSO), adaptive moment estimation with automatic differentiation (AD-Adam), and an inverse mapping solver network (IMS-Net) that directly infers the matching solution. Our main contributions are summarized as follows:
@@
  • Behavioral modeling of parasitic effects in TMNs via multiport network theory and DNN. First, we analyze the advantages of using scattering parameters (S-parameters) to characterize the circuit behavior of TMNs, and derive a general expression for the S-parameters of an arbitrary two-port TMN. Then, we explain the challenges associated with analytically constructing this general expression, and discuss the feasibility of employing deep learning for behavioral modeling. To characterize the circuit behavior of TMN while explicitly accounting for parasitic effects, we propose RECBM-Net, a DNN that learns the mapping between the TMN operating state and its corresponding S-parameters.
@@
  • Reformulating the data-driven adaptive impedance matching as a mathematical optimization problem. To address the challenges of formulating and solving conjugate matching equations due to the black-box nature of RECBM-Net in characterizing circuit behavior, we reformulate the adaptive impedance matching as a mathematical optimization problem based on the S-parameters. By varying the input to the RECBM-Net, the physical adjustment of tunable elements can be virtually emulated. The objective is to identify a TMN configuration whose S-parameters closely approximate the target transmission characteristics under the matched condition.
@@
  • Efficient matching solution determination strategies based on the RECBM-Net surrogate model. While brute-force grid search can identify the matching configuration, it incurs substantial online computational overhead due to repeated RECBM-Net evaluations. To reduce the cost of solving the previously formulated optimization problem, we first propose SAPSO, a hybrid algorithm that integrates global exploration with rapid convergence. In addition, we leverage automatic differentiation and employ the adaptive moment estimation (Adam) to enable efficient gradient-based search for a matching solution, further reducing computational overhead. To completely eliminate the need for repeated RECBM-Net evaluations, we introduce IMS-Net, a neural solver trained to directly predict the optimal matching solution.
@@
The remainder of this paper is organized as follows. Section II introduces the adaptive impedance matching system and formulates the matching problem while accounting for parasitic effects. In Section III, we propose a deep learning model RECBM-Net, designed to fit the circuit behavior of the TMN. Section IV presents three strategies for determining matching solutions based on the surrogate model of a practical TMN, each with varying computational overhead. The matching performance on a practical L-network is evaluated in Section V. Finally, Section VI concludes the paper.
@@ II System Model
In this section, we will begin by introducing the adaptive impedance matching system. Next, we will analyze the matching solution for the ideal L-network. Finally, we will examine in detail the complexities introduced by parasitic effects and formulate an adaptive impedance matching problem considering parasitic effects.
@@ II-A Adaptive Impedance Matching System
Consider the adaptive impedance matching system depicted in Fig. 1. It is composed of three main modules: a TMN, an impedance sensor, and a control unit. The TMN is typically composed of tunable capacitors and tunable inductors, with common configurations including L-network and $\pi$-network. It provides the desired impedance transformation by adjusting the values of reactive components. The impedance sensor detects impedance variations by analyzing the input and reflected signals extracted via a directional coupler. It can provide the control unit with key parameters such as the reflection coefficient and voltage standing wave ratio (VSWR). The control unit employs an adaptive impedance matching method to formulate a strategy for adjusting the TMN.
The control unit generally formulates the corresponding matching strategy based on two distinct types of measurement data [33]. One is the data that only indicates the degree of impedance mismatch, such as reflected power, VSWR, etc. Based on this type of data, the control unit can only adopt a trial-and-error impedance matching strategy. Impedance matching is achieved as quickly as possible by deciding the TMN configuration that is most worth trying each time. The other is the data that fully characterizes the complex impedance (e.g., reflection coefficient). Based on this type of data, the control unit knows the complex input impedance and can directly compute the matching solution based on the circuit model. In order to avoid the impracticality of frequent TMN adjustments, we focus on the latter scenario, in which the control unit can obtain the complex impedance. Also, we assume that the reflection coefficient is measured accurately.
@@ II-B Impedance Matching in Ideal L-network
Among the various TMN configurations, the L-network is particularly appealing for its simplicity and efficiency. In the ideal L-network depicted in Fig. 2, the tunable capacitors $C_{p}$ and $C_{s}$ enable the necessary impedance transformation. The input impedance $Z_{in}$ after impedance transformation by the L-network is given by
@@
where $B_{p}=\omega C_{p}$ and $B_{s}=\omega C_{s}$ are the susceptance of $C_{p}$ and $C_{s}$, respectively, $Z_{L}$ donates the load impedance and $\omega$ represents the angular frequency.
Based on Eq. (1), the reflection coefficient characterizes the input impedance can be given by
@@
The maximum power transfer theorem serves as the fundamental principle for tuning TMN to achieve impedance matching for different $\Gamma_{L}$. Maximum power is delivered when the input impedance is the complex conjugate of the source impedance. Since $R_{S}$ is typically a purely resistive 50 $\Omega$, this condition is reduced to $Z_{\text{in}}=R_{S}$, which in turn implies that the reflection coefficient $\Gamma_{\text{in}}$ measured by the impedance sensor is zero. Accordingly, the values of the $C_{p}$ and $C_{s}$ are adjusted to meet this requirement. Donating the load impedance as $Z_{L}=R_{L}+jX_{L}$ and substituting it into Eq. (1), we have the conjugate matching equations as
[…]
By solving Eq. (4), we have the closed-form expressions of the capacitor values required for impedance matching as
[…]

### v2

# A Data-Driven Adaptive Impedance Matching Method Robust to Parasitic Effects

@@ Abstract
Adaptive impedance matching between antennas and radio frequency front-end (RFFE) power modules is essential for mobile communication systems. To address the matching performance degradation caused by parasitic effects in practical tunable matching networks (TMNs), this paper proposes a data-driven adaptive impedance matching method that avoids physical adjustment. First, we propose the residual enhanced circuit behavior modeling network (RECBM-Net), a deep learning model that maps TMN operating states to their scattering parameters (S-parameters). Then, we formulate the matching process based on the trained surrogate model as a mathematical optimization problem. We employ two classic numerical methods with different online computational overhead, namely simulated annealing particle swarm optimization (SAPSO) and adaptive moment estimation with automatic differentiation (AD-Adam), to search for the matching solution. To further reduce the online inference overhead caused by repeated forward propagation through RECBM-Net, we train an inverse mapping solver network (IMS-Net) to directly predict the optimal solution. Simulation results show that RECBM-Net accurately predicts S-parameters, achieving a mean absolute error of $6.98\times 10^{-5}$. Across 9000 mismatched scenarios, the compliance rate after tuning increases from 0.97% with the analytical solution of the ideal L-network to 95.92% with SAPSO, 93.42% with AD-Adam, and 95% with IMS-Net. While AD-Adam significantly reduces computational overhead, lowering the average number of RECBM-Net inferences from 2097 with SAPSO to 285, it sacrifices some accuracy. IMS-Net requires only a single inference to obtain the matching solution, resulting in minimal online overhead while maintaining excellent matching accuracy.
@@ I Introduction
The impedance mismatch between the antenna and the radio frequency (RF) front-end power module reduces the power transmitted to the antenna and degrades overall system performance in mobile communication systems. Furthermore, impedance mismatch adversely affects the linearity of a digital predistortion power amplifier [1], resulting in signal modulation distortion and spectrum expansion. More critically, the reflected power caused by the mismatch may damage the RF front-end (RFFE) components [2]. Therefore, achieving impedance matching is crucial for mobile devices to maintain reliable and high-performance communication.
@@
The antenna impedance of mobile devices is influenced by multiple factors, e.g. operating frequency [3], various electromagnetic interactions arising from user holding methods [4, 5], user proximity [6, 7], and even user age and clothing [8]. The dynamic operating conditions of these devices induce impedance mismatches in the time-frequency domain. To mitigate the detrimental effects of impedance mismatch, extensive research has been conducted on adaptive impedance matching techniques for mobile communication systems.
@@
Conventionally, adaptive impedance matching is primarily achieved through numerical methods. The numerical methods use the feedback signal reflecting the degree of mismatch to determine the optimal matching solution through extensive trial and error. The works in [9, 10, 11] adopted sequential search strategies based on matching performance feedback. To enhance search efficiency, the authors in [12] employed various gradient-free optimization methods to minimize the magnitude of the reflection coefficient, such as the single-step and powell algorithms. Additionally, the work in [13, 14, 15] employed heuristic algorithms, while the work in [16] utilized gradient information to perform steepest gradient descent, reducing the number of trial-and-error iterations. The authors in [17] observed that when the parallel inductance of the T-network is fixed, the real part of the input impedance follows a linear fractional relationship with the reactance of the first series element, and proposed a binary search algorithm to reduce the number of iterations. The disadvantages of numerical methods are their inefficiency and riskiness. As the matching network components must be physically adjusted during each trial, this tuning process introduces both amplitude and phase modulation to the radiated signal, which can lead to data corruption if tuning occurs during transmission [18]. Therefore, numerical methods prove impractical for achieving adaptive impedance matching in mobile devices.
@@
Compared to numerical methods, analytical methods derive the matching solution directly from the circuit model and accurate complex impedance information, thereby avoiding the iterative trial-and-error procedures. The authors in [19] proposed a method for real-time measurement of the complex impedance at any location within the RFFE of a handset based on perturbation theory, providing critical support for the analytical computation of the matching solution. The work of [20] presented explicit mathematical expressions for the values of components when the ideal $\pi$-network is perfectly matched. Furthermore, the authors in [21] developed an orthogonal detection technique that directly sets the real and imaginary parts of the impedance. To address imperfect matching caused by the limited tuning range of practical $\pi$-network, the work of [22] proposed an analytical method to compute the optimal solution within the tunable range. Additionally, the work of [23] considered the parasitic effects introduced by microelectromechanical systems (MEMS) components in the practical $\pi$-network and provided an analytical solution by solving the conjugate matching equation of the equivalent circuit, which remains an approximation of the practical circuit. As analytical methods are inherently model-dependent, the accuracy of the matching is determined by the degree of conformity between the practical circuit model and the model assumed by the algorithm.
@@
Recently, powerful artificial intelligence (AI) technologies have been deployed to realize efficient and accurate adaptive impedance matching. To address impedance mismatch caused by frequency changes, the work in [24] proposed a deep learning model that maps a target frequency response of the reflection coefficient to the corresponding L-network configuration for a given planar inverted-F antenna. Considering the complexity of the model, the authors in [25] proposed a low-complexity shallow learning model based on ridge regression to achieve the same mapping. In addition, to address the impedance mismatch caused by variations in the distance between the transmitting and receiving coils in wireless power transfer systems, the work in [26] employed a neural network to establish a mapping between the equivalent load impedance and the corresponding matching solution at a specific frequency. Furthermore, considering the time-frequency domain mismatch, the work of [27] proposed a deep neural network (DNN) to directly determine the ideal $\pi$-network matching solution based on incomplete impedance information. The construction of large and accurate supervised datasets is critical to the success of AI-based adaptive impedance matching methods.
@@
The above analytical and AI-based methods are applicable only to tunable matching networks (TMNs) that either neglect parasitic effects or consider only simplified cases. However, parasitic effects in practical TMNs are unavoidable and complex due to inherent limitations in the process of manufacturing and packaging of RF circuits. For example, common structures such as traces, bonding wires [28], vias [29], pins [30], solder balls [31], etc., will introduce parasitic capacitance, inductance, and resistance. Particularly in mobile communication systems, TMNs are often implemented using discrete electronically switchable inductor and capacitor banks [32] to satisfy miniaturization constraints. The dense integration of switches, transmission lines, and non-ideal discrete components significantly aggravates parasitic effects.
@@
Parasitic effects render the equivalent circuit topology of a practical TMN unknown. When these effects are taken into account, matching solutions derived from analytical methods based on idealized or specific topologies often fail to meet the actual matching requirements. Similarly, for the aforementioned AI-based methods, datasets constructed from idealized topologies fail to reflect actual matching behavior, while those based on practical topologies suffer from the difficulty of acquiring accurate labels. These issues lead to degraded matching performance in AI-based methods as well. To the best of our knowledge, achieving efficient and accurate adaptive impedance matching remains an open problem when parasitic effects cause the practical TMN to deviate from its idealized topology.
@@
To address this challenge, in this paper, we propose a data-driven adaptive impedance matching method. First, we analyze the impact of parasitic effects on the behavior of a TMN using multiport network theory. Considering the high nonlinearity and uncertainty introduced by these parasitic effects, we propose a deep learning model, residual enhanced circuit behavior modeling network (RECBM-Net), to accurately learn the system behavior. Based on the trained RECBM-Net, we further develop three matching solution determination strategies with varying online computational overhead: simulated annealing particle swarm optimization (SAPSO), adaptive moment estimation with automatic differentiation (AD-Adam), and an inverse mapping solver network (IMS-Net) that directly infers the matching solution. Our main contributions are summarized as follows:
@@
  • Behavioral modeling of parasitic effects in TMNs via multiport network theory and the DNN. First, we analyze the advantages of using scattering parameters (S-parameters) to characterize the circuit behavior of TMNs and derive a general expression for the S-parameters of an arbitrary two-port TMN. Then, we explain the challenges associated with analytically constructing this general expression and discuss the feasibility of employing deep learning for behavioral modeling. To characterize the circuit behavior of a TMN while explicitly accounting for parasitic effects, we propose RECBM-Net, a DNN that learns the mapping between the TMN’s operating state and its corresponding S-parameters.
@@
  • Reformulating data-driven adaptive impedance matching as a mathematical optimization problem. To address the challenges of formulating and solving conjugate matching equations due to the black-box nature of RECBM-Net in characterizing circuit behavior, we reformulate the adaptive impedance matching as a mathematical optimization problem based on the S-parameters. By varying the input to RECBM-Net, the physical adjustment of tunable elements can be virtually emulated. The objective is to identify a TMN configuration whose S-parameters closely approximate the target transmission characteristics under the matched condition.
@@
  • Efficient matching solution determination strategies based on the RECBM-Net surrogate model. While brute-force grid search can identify the matching network configuration, it incurs substantial online computational overhead due to repeated RECBM-Net evaluations. To reduce the cost of solving the previously formulated optimization problem, we first propose SAPSO, a hybrid algorithm that integrates global exploration with rapid convergence. In addition, we leverage automatic differentiation (AD) and employ the adaptive moment estimation (Adam) to enable efficient gradient-based search for a matching solution, further reducing computational overhead. To completely eliminate the need for repeated RECBM-Net evaluations, we introduce IMS-Net, a neural solver trained to directly predict the optimal matching solution.
@@
The remainder of this paper is organized as follows. Section II introduces an adaptive impedance matching system and formulates the matching problem while accounting for parasitic effects. In Section III, we propose a deep learning model, RECBM-Net, designed to fit the circuit behavior of the TMN. Section IV presents three strategies for determining matching solutions based on the surrogate model of a practical TMN, each with varying computational overhead. The matching performance on a practical L-network is evaluated in Section V. Finally, Section VI concludes the paper.
@@ II System Model
In this section, we will begin by introducing the adaptive impedance matching system. Next, we will analyze the matching solution for the ideal L-network. Finally, we will examine in detail the complexities introduced by parasitic effects and formulate an adaptive impedance matching problem that accounts for parasitic effects.
@@ II-A Adaptive Impedance Matching System
Consider the adaptive impedance matching system depicted in Fig. 1. It is composed of three main modules: a TMN, an impedance sensor, and a control unit. The TMN is typically composed of tunable capacitors and tunable inductors, with common configurations including the L-network and the $\pi$-network. It provides the desired impedance transformation by adjusting the values of the reactive components. The impedance sensor detects impedance variations by analyzing the incident and reflected signals extracted via a directional coupler. It can provide the control unit with key parameters such as the reflection coefficient and voltage standing wave ratio (VSWR). The control unit employs an adaptive impedance matching method to formulate a strategy for adjusting the TMN.
The control unit generally formulates the corresponding matching strategy based on two distinct types of measurement data [33]. One is the data that only indicates the degree of impedance mismatch, such as reflected power and VSWR. Based on this type of data, the control unit can only adopt a trial-and-error impedance matching strategy. Impedance matching is achieved as quickly as possible by deciding the TMN configuration that is most worth trying each time. The other is the data that fully characterizes the complex impedance (e.g., reflection coefficient). Based on this type of data, the control unit knows the complex input impedance and can directly compute the matching solution based on the circuit model. In order to avoid the impracticality of frequent TMN adjustments, we focus on the latter scenario, in which the control unit can obtain the complex impedance. Also, we assume that the reflection coefficient is measured accurately.
@@ II-B Impedance Matching in Ideal L-network
Among the various TMN configurations, the L-network is particularly appealing for its simplicity and efficiency. In the ideal L-network depicted in Fig. 2, the tunable capacitors $C_{p}$ and $C_{s}$ enable the necessary impedance transformation. The input impedance $Z_{\text{in}}$ after impedance transformation by the L-network is given by
@@
where $B_{p}=\omega C_{p}$ and $B_{s}=\omega C_{s}$ are the susceptances of $C_{p}$ and $C_{s}$, respectively, $Z_{L}$ denotes the load impedance and $\omega$ represents the angular frequency.
Based on Eq. (1), the reflection coefficient that characterizes the input impedance can be given by
@@
The maximum power transfer theorem serves as the fundamental principle for tuning a TMN to achieve impedance matching for different $\Gamma_{L}$. Maximum power is delivered when the input impedance is the complex conjugate of the source impedance. Since $R_{S}$ is typically a purely resistive 50 $\Omega$, this condition is reduced to $Z_{\text{in}}=R_{S}$, which in turn implies that the reflection coefficient $\Gamma_{\text{in}}$ measured by the impedance sensor is zero. Accordingly, the values of the $C_{p}$ and $C_{s}$ are adjusted to meet this requirement. Denoting the load impedance as $Z_{L}=R_{L}+jX_{L}$ and substituting it into Eq. (1), we write the conjugate matching equations as
$\left\{\begin{aligned} &\frac{B_{s}^{2}R_{L}}{(B_{p}B_{s}R_{L})^{2}+(B_{p}+B_{s}-B_{p}B_{s}X_{L})^{2}}=R_{S},\\ &\frac{2B_{p}B_{s}X_{L}+B_{s}^{2}X_{L}-B_{p}B_{s}^{2}(R_{L}^{2}+X_{L}^{2})-B_{p}-B_{s}}{(B_{p}B_{s}R_{L})^{2}+(B_{p}+B_{s}-B_{p}B_{s}X_{L})^{2}}=0.\end{aligned}\right$. (4)
By solving Eq. (4), we obtain the closed-form expressions of the capacitor values required for impedance matching as
$\left\{\begin{aligned} &C_{p}=\frac{R_{L}R_{S}-R_{L}^{2}\pm X_{L}\sqrt{R_{L}R_{S}-R_{L}^{2}}}{\omega[R_{L}X_{L}R_{S}\pm R_{L}R_{S}\sqrt{R_{L}R_{S}-R_{L}^{2}}]},\\ &C_{s}=\frac{X_{L}\pm\sqrt{R_{L}R_{S}-R_{L}^{2}}}{\omega[R_{L}^{2}+X_{L}^{2}-R_{L}R_{S}]}.\end{aligned}\right$. (5)